# Mitigating Feature Gap for Adversarial Robustness by Feature Disentanglement

### v1

@@ Abstract
Deep neural networks are vulnerable to adversarial samples. Adversarial fine-tuning methods aim to enhance adversarial robustness through fine-tuning the naturally pre-trained model in an adversarial training manner. However, we identify that some latent features of adversarial samples are confused by adversarial perturbation and lead to an unexpectedly increasing gap between features in the last hidden layer of natural and adversarial samples. To address this issue, we propose a disentanglement-based approach to explicitly model and further remove the latent features that cause the feature gap. Specifically, we introduce a feature disentangler to separate out the latent features from the features of the adversarial samples, thereby boosting robustness by eliminating the latent features. Besides, we align features in the pre-trained model with features of adversarial samples in the fine-tuned model, to further benefit from the features from natural samples without confusion. Empirical evaluations on three benchmark datasets demonstrate that our approach surpasses existing adversarial fine-tuning methods and adversarial training baselines.
Mitigating Feature Gap, Feature Disentanglement, Adversarial Robustness, Adversarial Fine-Tuning
Figure 1: Toy illustration of Disentanglement. We model features of adversarial samples and then remove the latent features confused by adversarial perturbation to correct the prediction.
1 Introduction
@@
Deep neural networks (DNNs) have shown impressive performances in various domains of machine learning. However, it has been demonstrated that DNNs are susceptible to adversarial samples and the prediction can be easily manipulated (Goodfellow et al., 2014). Adversarial samples deceive DNNs by introducing imperceptible noise to clean samples. With the widespread use of DNNs, the presence of adversarial samples poses a growing potential threat, making it crucial to enhance the robustness of networks.
@@
Many defensive techniques are proposed including adversarial training (AT) (Madry et al., 2017; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020; Huang et al., 2020; Dong et al., 2021; Jia et al., 2022; Jin et al., 2022; Zhou et al., 2022), adversarial fine-tuning (AFT) (Suzuki et al., 2023; Maini et al., 2020; Madaan et al., 2021; Croce & Hein, 2022; Jeddi et al., 2020; Zhu et al., 2023; Si et al., 2020; Agarwal et al., 2023; Liu et al., 2023), adversarial purification (Yoon et al., 2021; Sun et al., 2023; Lee & Kim, 2023; Allen-Zhu & Li, 2022; Nie et al., 2022), adversarial detection (Hickling et al., 2023; Zheng & Hong, 2018; Pang et al., 2018), etc. It is universally acknowledged that AT is the most effective defensive technique, but it suffers from several times the training cost as standard (natural) training. To save the cost of training time, AFT employs the same loss function as AT to fine-tune the pre-trained model within a few epochs.
@@
We identify an issue of the increasing feature gap by analyzing three models fine-tuned by three AFT methods (vanilla AFT, AFT+RGKD, ARREST) (Suzuki et al., 2023; Jeddi et al., 2020). As illustrated in Figure 2, the feature distance between natural and adversarial samples abnormally maintains an increasing direction during the fine-tuning process. Though ARREST initially exhibits notable fluctuations with a decreasing trend, the overall magnitudes of feature distance in three AFT methods show an increasing trend. We suppose that some latent features are easy to be confused by adversarial attacks and lead to the growing gap between the features of natural and adversarial samples. Considering that a robust model should treat natural and adversarial samples equally and extract similar features from them, we expect to remove the latent features causing the feature gap to achieve better robust performance.
@@
In this paper, we analyze the latent features causing the feature gap and introduce an approach based on disentanglement to explicitly model and remove these features. We propose a new technique named Adversarial Fine-tuning via Disentanglement (AFD). First, we disentangle the features of adversarial samples as two components, in which the latent features causing the feature gap are identified as the features confused by adversarial perturbation. Next, we propose a feature disentangler to model and separate them from the features of adversarial samples. We maximize the predicted probabilities of the wrong predicted classes to acquire the latent features. Then we impose a constraint to distance the features of adversarial samples from them, to eliminate the latent features and mitigate the feature gap. Besides, we align features of adversarial samples in the fine-tuned model with features of natural samples in the naturally pre-trained model, further eliminating the latent features. Experiments demonstrate that our AFD alleviates the gap between features of natural and adversarial samples and improves the robustness. The key contributions are summarized as follows:
@@
  • We observe the gap in features between natural and adversarial samples anomalously increases when applying adversarial fine-tuning methods, resulting from the latent features confused by adversarial perturbation.
@@
  • We propose a disentanglement-based approach, which explicitly models and further eliminates the latent features causing the feature gap by a disentangler and loss function of alignment.
@@
2 Related Work
@@
Adversarial Training Adversarial training (AT) is the most effective technique to defend against attacks. Numerous studies have been conducted to boost robustness and improve the trade-off between accuracy and robustness. Madry et al. (2017) propose Projected Gradient Descent (PGD) attack and PGD-based adversarial training (PGD-AT), compelling the model to correctly classify adversarial samples within the epsilon sphere during training, which is the pioneer in adversarial learning. Zhang et al. (2019) reduce the divergence of probability distributions of natural and adversarial samples to mitigate the difference between robust and natural accuracy. Wang et al. (2020) find that misclassified samples harm adversarial robustness significantly, and propose to improve the model’s attention to misclassification by adaptive weights. Huang et al. (2020) replace labels with soft labels predicted by the model and adaptively reduce the weights of misclassified samples to alleviate the robust overfitting problem. Dong et al. (2021) also propose a similar idea of softening labels and explain the different effects of hard and soft labels on robustness by investigating the memory behavior of the model for random noisy labels. Zhou et al. (2022) embed a label transition matrix into models to infer natural labels from adversarial noise. Though AT has good robust performances, it suffers from a large quantity of computing expenses and training costs. Some researchers have developed approaches to accelerate adversarial training, commonly known as fast adversarial training (FAT) (Wong et al., 2020; Kim et al., 2020; Huang et al., 2023). Since these FAT methods only require a little training time like AFT, we compare them with our method.
@@
Adversarial Fine-tuning Adversarial fine-tuning (AFT) has been employed to enhance the pre-trained model within a few epochs for various targets, particularly for achieving adversarial robustness. Moosavi-Dezfooli et al. (2018) analyze the effect of AFT by comparing the decision boundary of a DNN before and after applying AFT. Some researchers (Tramer & Boneh, 2019; Maini et al., 2020; Madaan et al., 2021; Croce & Hein, 2022) aim to achieve multiple $l_{p}$ adversarial robustness through specific strategies. Jeddi et al. (2020) employ vanilla AFT to expedite the training of robust models with a few training epochs and a warm-up learning rate schedule. Based on representation learning, Suzuki et al. (2023) propose ARREST and AFT+RGKD, which reduce the feature gap between natural and adversarial samples. Zhu et al. (2023) propose a metric to measure the robustness of modules and fine-tune them for better out-of-distribution performance. Liu et al. (2023) introduce a two-pipeline structure to improve the relationship between the weight norm and its gradient norm in batch normalization layers. Despite the extensive work, few methods (vanilla AFT, ARREST, and AFT+RGKD) are based on naturally pre-trained models, which are important baselines in our paper.
3 Methodology
In this section, we first reveal the issue of the increasing gap of features between natural and adversarial samples in AFT. Then we introduce a disentanglement-based method, to model and further remove the latent features leading to the feature gap. Our method is expected to mitigate the feature gap to enhance adversarial robustness.
3.1 Preliminaries
@@
We denote a DNN as $h(x)=\omega\cdot g(x;\theta_{g})$, where $x$ is the input image, and $\theta_{g}$ is the parameters of the feature extractor $g$, i.e., the model before the last linear classifier. $\omega$ denotes the parameters of the last linear classifier. The parameters of the naturally pre-trained and fine-tuned model are denoted as $\theta^{P}$ and $\theta^{F}$, respectively. The architecture and dataset for both pre-training and fine-tuning remain consistent, and the model parameters $\theta^{F}$ are initialized by $\theta^{P}$. $x$ denotes the clean sample, and $x^{\prime}$ denotes the adversarial sample. ${\|{\delta}\|}_{p}\leq\epsilon$ denotes the adversarial perturbation, where $\epsilon$ represents the budget of adversarial attack. We use ${\|{\cdot}\|}_{p}$ to denote the $l_{p}$ norm. Adversarial and natural features denote the features in the last hidden layer of adversarial and natural samples, i.e., $g(x^{\prime};\theta_{g})$ and $g(x;\theta_{g})$ respectively. $\mathcal{L}$ denotes the cross-entropy loss.
3.2 Model the Latent Feature causing Feature Gap
@@
We evaluate the feature gap in AFT by measuring the ${L}_{p}$ distance between natural and adversarial features. As shown in Figure 2, the overall magnitudes of feature distance between natural and adversarial samples in three AFT methods (vanilla AFT, ARREST, AFT+RGKD) are becoming larger. The gap between the two types of features in ARREST initially decreases and then increases, which exhibits a V-shape trend. In addition, the feature gap in other AFT methods presents a consistently increasing trend. Despite AFT enhancing robust accuracy during training, the feature distance in AFT maintains a growing trend overall. Intuitively, the ideal robust model should have a consistent understanding and analysis of natural and adversarial samples, thus extracting similar (the same) features from them. Figure 2 indicates that existing AFT methods neglect the gap of features between natural and adversarial samples. Narrowing the feature gap is expected to achieve better performance and enhance robust accuracy.
@@
To realize the target, we can disentangle two types of features from the features of natural and adversarial samples. We define natural features $f_{nat}$ as the unconfused features $f_{u}$. We can define adversarial features $f_{adv}$ as the combination of the unconfused features $f_{u}$ and the confused features $f_{c}$, the latter of which leads to the gap in features between natural and adversarial samples. In this modeling, the unconfused features represent the features that contribute to accurate classification, which are easily extracted from natural samples. Besides, the confused features are identified as the latent features causing the gap between features of natural and adversarial samples, which are confused by the adversarial perturbation. As illustrated in Figure 1, if the confused features are removed in adversarial features, the features of adversarial samples can align with those of natural samples, resulting in a zero feature gap and accurate predictions for adversarial samples.
Figure 2: $L_{p}$ distances of features between natural and adversarial samples on ResNet18 on CIFAR-10. The horizontal axis denotes the training epoch. Despite the initial decreasing fluctuations observed in ARREST and AFT+RGKD, AFT methods exhibit an increasing trend on the whole.
@@
We suppose the phenomenon in Figure 2 is attributable to the presence of the confused features $f_{c}$. Because the confused features persist as latent features and don’t diminish during fine-tuning, the gap between features of natural and adversarial samples cannot be bridged. The confused features induce the gap between natural and adversarial features, leading to the limited robust performance. Based on our observations and analyses, we introduce a new approach named Adversarial Fine-tuning via Disentanglement (AFD), as shown in Figure 3. It applies a feature disentangler to explicitly extract the confused features, and remove them by the elimination and alignment of features.
Figure 3: Overview of our AFD. Lock denotes the frozen parameters. The disentangler disentangles adversarial features into $f_{1}$ and $f_{2}$. We obtain confused features $f_{2}$ by matching the wrong predicted labels in $\mathcal{L}_{1}$. Next, we employ $\mathcal{L}_{2}$ to eliminate confused features from unconfused features $f_{1}$. Besides, we apply $\mathcal{L}_{3}$ to further align $f_{1}$ with pre-trained unconfused features $f_{nat}^{\prime}$ to correct the prediction.
3.3 Adversarial Fine-tuning Based on Disentanglement
@@
The primary strategy is to explicitly disentangle the unconfused and confused features from the features of adversarial samples. We propose a feature disentangler to model confused features for further elimination. The features in the last hidden layer are input into the disentangler, yielding two disentangled feature vectors. We define these feature vectors as 1) $f_{1}$, features contributing to correct classification without confusion; and 2) $f_{2}$, features inducing wrong predictions of adversarial samples. To implement this disentangling, we introduce a disentangler with two linear blocks $D_{1},D_{2}$ between the feature extractor $g$ and the linear classifier $\omega$. These linear blocks are expected to disentangle adversarial features into the unconfused features and confused features. The disentangled feature vectors $f_{1}$ and $f_{2}$ are constrained to match the unconfused features and confused features, formulated as:
@@
$\begin{split}f_{1}&=\ {D_{1}}(g(x^{\prime};\theta_{g}^{F});\theta_{D_{1}}),\\ f_{2}&=\ {D_{2}}(g(x^{\prime};\theta_{g}^{F});\theta_{D_{2}}),\\ {\rm where}\ &\left\{\begin{array}[]{lr}f_{1}\stackrel{{\scriptstyle D_{1}}}{{\longrightarrow}}f_{u},&\\ f_{2}\stackrel{{\scriptstyle D_{2}}}{{\longrightarrow}}f_{c},&\end{array}\right.\end{split}$ (1)
@@
where $\theta_{g}^{F}$ denotes the parameters of the feature extractor in the fine-tuned model. We take $f_{1}$ to correctly classify samples and $f_{2}$ to match the confused features. The input and output dimensions of the linear block $D_{1},D_{2}$ are set to match the channel dimensions of the feature vector of adversarial samples before disentanglement. This ensures the parameters of the linear block $\theta_{D_{1}}$ can be seamlessly integrated with the parameter of the last linear classifier $\omega^{F}$, expressed as $\theta_{D_{1}}\cdot\omega^{F}={\omega^{F\prime}}$. Thus we only need to load the linear classifier $\omega^{F\prime}$ without additional disentangling modules in the deployment (test) phase.
@@
To achieve good performance on both accuracy and robustness, we should design suitable loss functions to realize Formula 1. First, we suggest fitting $f_{2}$ to incorrectly predicted labels $y^{\prime}$ to model the confused features. The confused features result in obvious shifts in both features and predicted probabilities, leading to wrong predicted class labels in most cases. Thus, it can be identified as the latent features causing the gap between natural and adversarial features. To disentangle the confused features from adversarial features, we force the disentangled features $f_{2}$ to match incorrect class labels, formulated as:
@@
$\begin{split}\mathcal{L}_{1}=\ &{\rm\mathop{arg\ min}\limits_{\theta_{D_{2}}}}\ \mathcal{L}(\omega^{F}\cdot f_{2},y^{\prime}),\\ {\rm where}\quad&y^{\prime}={\rm\mathop{arg\ sort}\limits_{y^{\prime}\neq y}}\ \omega^{F}\cdot f_{1},\end{split}$ (2)
@@
where $y^{\prime}$ denotes wrong predicted labels with the maximum predicted probability, and sort represents finding the index of the maximum value. We only optimize the disentangling block $D_{2}$ instead of the entire model. This selective optimization prevents potential harm to the feature extractor that could result from learning the confused features.
@@
In addition, we introduce a constraint on $f_{1}$ to keep away from the confused features $f_{2}$, removing the confused features from the features of adversarial samples. The exclusive force between $f_{1}$ and $f_{2}$ encourages the elimination of the confused features from the features of adversarial samples. This constraint aims to protect adversarial features from being perturbed by the confused features, and then gain the unconfused features $f_{1}$. Given by:
$\mathcal{L}_{2}={\rm\mathop{arg\ max}\limits_{\theta^{F},\theta_{D_{1}}}}\ \mathcal{D}(f_{1},\bar{f_{2}}),$ (3)
@@
where $\mathcal{D}$ denotes a distance function, $\bar{f_{2}}$ means the gradient back-propagation of $f_{2}$ is frozen.
3.4 Adversarial Fine-tuning Based on Alignment
@@
The second strategy is to align the adversarial features in the fine-tuned model with the natural features in the pre-trained model. Since the pre-trained model has a high natural accuracy and doesn’t suffer from adversarial attacks, the natural features in the pre-trained model can be identified as ideal unconfused features. Thus, adversarial features in the fine-tuned model can be greatly improved by being aligned with the natural features in the pre-trained model. Besides, lots of work (Zhang et al., 2019; Wang et al., 2020; Suzuki et al., 2023) conducts an alignment between natural and adversarial features, which is regarded to mitigate the trade-off between accuracy and robustness. We take the disentangled features $f_{1}$ to match the natural features in the pre-trained model, further removing the confused features from $f_{1}$. The AFT loss function with the alignment can be formulated as:
$\mathcal{L}_{3}={\rm\mathop{arg\ min}\limits_{\theta^{F},\theta_{D_{1}}}}\ \mathcal{L}(\omega^{F}\cdot f_{1},y)+\gamma\cdot\mathcal{D}(f_{1},g(x;\theta_{g}^{P})),$ (4)
@@
where $\theta_{g}^{P}$ denotes parameters of the feature extractor in the pre-trained model, and $\gamma$ denotes the weight of the alignment. Therefore, the objective function of our AFD is defined as follows:
@@
where $\alpha$ and $\beta$ denote the weights of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The algorithm of AFD is as follows:

### v2

@@ Abstract
Adversarial fine-tuning methods enhance adversarial robustness via fine-tuning the pre-trained model in an adversarial training manner. However, we identify that some specific latent features of adversarial samples are confused by adversarial perturbation and lead to an unexpectedly increasing gap between features in the last hidden layer of natural and adversarial samples. To address this issue, we propose a disentanglement-based approach to explicitly model and further remove the specific latent features. We introduce a feature disentangler to separate out the specific latent features from the features of the adversarial samples, thereby boosting robustness by eliminating the specific latent features. Besides, we align clean features in the pre-trained model with features of adversarial samples in the fine-tuned model, to benefit from the intrinsic features of natural samples. Empirical evaluations on three benchmark datasets demonstrate that our approach surpasses existing adversarial fine-tuning methods and adversarial training baselines.
Introduction
@@
Deep neural networks (DNNs) have shown impressive performances in various domains of machine learning. However, it has been demonstrated that DNNs are susceptible to adversarial samples and the prediction can be easily manipulated (Goodfellow, Shlens, and Szegedy 2014). Adversarial samples deceive DNNs by introducing imperceptible noise to clean samples. The presence of adversarial samples poses a growing potential threat (Madry et al. 2017; Croce and Hein 2020; Yu, Gao, and Xu 2021; Wei et al. 2022, 2023; Liu et al. 2024; Xia et al. 2024), making it crucial to enhance the robustness of networks.
Figure 1: Illustration of our motivation. (a) $L_{\infty}$ distances of features between natural and adversarial samples during fine-tuning. Previous AFT methods exhibit an increasing trend of the feature gap. (b) Toy illustration of Disentanglement. We model features of adversarial samples and then remove the specific latent features confused by adversarial perturbation to correct the prediction.
@@
Many defensive techniques are proposed including adversarial training (AT) (Madry et al. 2017; Zhang et al. 2019; Wang et al. 2020; Wu, Xia, and Wang 2020; Huang, Zhang, and Zhang 2020; Dong et al. 2021; Jia et al. 2022; Jin et al. 2022; Yu et al. 2022; Zhou et al. 2022), adversarial fine-tuning (AFT) (Suzuki et al. 2023; Maini, Wong, and Kolter 2020; Madaan, Shin, and Hwang 2021; Croce and Hein 2022; Jeddi, Shafiee, and Wong 2020; Zhu et al. 2023; Si et al. 2020; Agarwal et al. 2023; Liu et al. 2023), adversarial purification (Yoon, Hwang, and Lee 2021; Sun et al. 2023; Lee and Kim 2023; Allen-Zhu and Li 2022; Nie et al. 2022), adversarial detection (Hickling, Aouf, and Spencer 2023; Zheng and Hong 2018; Pang et al. 2018), etc. AT is regarded as the most effective defensive technique among them, but it suffers from several times the training cost as standard (natural) training. To save the cost of training time, AFT employs the same loss function as AT to fine-tune the pre-trained model within a few epochs, which has aroused wide attention.
@@
We identify an issue of the increasing feature gap by analyzing three models fine-tuned by three AFT methods: vanilla AFT (Jeddi, Shafiee, and Wong 2020), AFT+RGKD, ARREST (Suzuki et al. 2023). As illustrated in Figure 1 (a), the feature distance between natural and adversarial samples abnormally maintains an increasing direction during the fine-tuning process. Though ARREST initially exhibits notable fluctuations with a decreasing trend, the overall magnitudes of feature distance in three AFT methods show an increasing trend. We suppose that some specific latent features are easily confused by adversarial attacks and lead to the growing gap between the features of natural and adversarial samples. Considering that a robust model should treat natural and adversarial samples equally and extract similar features from them, we expect to remove the specific latent features causing the feature gap to achieve better robust performance, as shown in Figure 1 (b).
@@
In this paper, we analyze the specific latent features causing the feature gap and introduce an approach based on disentanglement to explicitly model and remove these features. We propose a new technique named Adversarial Fine-tuning via Disentanglement (AFD). First, we disentangle the features of adversarial samples as two components, in which the specific latent features causing the feature gap are identified as the features confused by adversarial perturbation. Next, we propose a feature disentangler to model and separate them from the features of adversarial samples. We maximize the predicted probabilities of the wrong predicted classes to acquire the specific latent features. Then we impose a constraint to distance the features of adversarial samples from them, to eliminate the specific latent features and mitigate the feature gap. Besides, we align features of adversarial samples in the fine-tuned model with features of natural samples in the naturally pre-trained model, further eliminating the specific latent features. Experiments demonstrate that our AFD alleviates the feature gap and improves the robustness. Contributions are summarized as follows:
@@
  • We observe the gap in features between natural and adversarial samples anomalously increases in adversarial fine-tuning methods, resulting from the specific latent features confused by adversarial perturbation.
@@
  • We provide the theoretical analysis of feature disentanglement. Then we introduce a disentanglement-based AFT approach, which explicitly models and further eliminates the specific latent features causing the feature gap by disentanglement and alignment.
@@
Related Work
Adversarial Attack Adversarial Attack is first introduced in Goodfellow, Shlens, and Szegedy (2014), which fools DNNs by imperceptible perturbations. Madry et al. (2017) propose Projected Gradient Descent (PGD) attack based on the gradient projection and random start. Croce and Hein (2020) introduce a powerful adaptive attack (AutoAttack), including three white-box attacks and a black-box attack. Yu, Gao, and Xu (2021) demonstrate that exploiting latent features is highly effective against many defense techniques. Besides, adversarial attacks can also work in the physical world. Wei et al. (2022) present a comprehensive overview of physical adversarial attacks. Wei et al. (2023) propose a novel physical adversarial attack that applies the Warming Paste and Cooling Paste to hide persons from being detected. Liu et al. (2024) learn adversarial semantic perturbations in the latent space for high attack capabilities and low perceptibility. We take PGD and AutoAttack to evaluate robustness.
@@
Adversarial Training Adversarial training (AT) is the most effective technique to defend against attacks. Madry et al. (2017) propose PGD-based adversarial training (PGD-AT), compelling the model to correctly classify adversarial samples within the epsilon sphere during training. Zhang et al. (2019) reduce the divergence of probability distributions of natural and adversarial samples to mitigate the difference between robust and natural accuracy. Xie et al. (2019) apply non-local means or other filters as the denoising block. Wang et al. (2020) find that misclassified samples harm adversarial robustness significantly, and propose to improve the model’s attention to misclassification by adaptive weights. Yan et al. (2021) manipulates channels’ activations to align the channel with the relevance to predictions. Yang et al. (2021) introduce a disentanglement-based architecture to generate class-specific and class-irrelevant representations. Zhou et al. (2022) embed a label transition matrix into models to infer natural labels from adversarial noise. Kim et al. (2023) propose the FSR module to separate non-robust activations and recalibrate the features. Zhou et al. (2023) define two attributes of robust features to guide robust training. Though AT has good robust performances, it suffers from a large quantity of computing expenses. Some work has been developed to accelerate adversarial training, commonly known as fast adversarial training (FAT) (Wong, Rice, and Kolter 2020; Kim, Lee, and Lee 2020; Huang et al. 2023). Since these FAT methods only require a little training time like AFT, we compare them with our method.
@@
Adversarial Fine-tuning Adversarial fine-tuning (AFT) has been employed to enhance the pre-trained model within a few epochs for various targets, particularly for achieving adversarial robustness. Moosavi-Dezfooli et al. (2018) analyze the effect of AFT by comparing the decision boundary of a DNN before and after applying AFT. Kumari et al. (2019); Yu, Gao, and Xu (2021) discover the vulnerability of latent features in adversarial trained models and propose adversarial approaches to mitigate this. Jeddi, Shafiee, and Wong (2020) employ vanilla AFT to expedite the training of robust models with a few training epochs and a warm-up learning rate schedule. Some researchers (Tramer and Boneh 2019; Maini, Wong, and Kolter 2020; Madaan, Shin, and Hwang 2021; Croce and Hein 2022) aim to achieve multiple $l_{p}$ adversarial robustness through specific strategies. Based on representation learning, Suzuki et al. (2023) constrain the representation to mitigate the accuracy-robustness trade-off. Although they aim to learn suitable features, the gap between natural and adversarial features can’t be eliminated by simple feature constraint, which has been indicated in Figure 1 (a)Zhu et al. (2023) propose a metric to measure the robustness of modules and fine-tune them for better out-of-distribution performance. Liu et al. (2023) introduce a two-pipeline structure to improve the relationship between the weight norm and its gradient norm in batch normalization layers. Despite the extensive work, few methods (vanilla AFT (Jeddi, Shafiee, and Wong 2020), ARREST and AFT+RGKD (Suzuki et al. 2023)) are based on naturally pre-trained models, which are our important baselines.
Figure 2: Overview of our AFD. Lock denotes the frozen parameters, $BZ$ denotes the batch size, $Cls$ denotes the class number, and numbers in brackets denote the sizes of features or parameters. During fine-tuning, Adversarial features are disentangled into $f_{1}$ and $f_{2}$. $\mathcal{L}_{CE}$ ensures $f_{2}$ to approximate confused features. We keep $f_{1}$ away from $f_{2}$ to eliminate confused features in $f_{1}$. Besides, we align $f_{1}$ with pre-trained intrinsic features $f_{nat}^{\prime}$ to further correct the prediction. During inference, we multiply $D_{1}$ and ${\omega}^{F}$ to obtain the robust model without additional modules.
Preliminaries
@@
We denote a DNN as $h(x)=\omega\cdot g(x;\theta_{g})$, where $x$ is the input image, and $\theta_{g}$ is the parameters of the feature extractor $g$, i.e., the model before the last linear classifier. $\omega$ denotes the parameters of the last linear classifier. The parameters of the naturally pre-trained and fine-tuned model are denoted as $\theta^{P}$ and $\theta^{F}$, respectively. The model parameters $\theta^{F}$ are initialized by $\theta^{P}$. $x$ denotes the clean sample, and $x^{\prime}$ denotes the adversarial sample. $\mathbb{N}(x,\epsilon)$ represents $\{x^{\prime}:\|x^{\prime}-x\|_{\infty}\leq\epsilon\}$, $\epsilon$ is the perturbation budget, ${\|{\cdot}\|}_{p}$ denotes the $l_{p}$ norm. Adversarial and natural features denote the features in the last hidden layer of adversarial and natural samples, i.e., $g(x^{\prime};\theta_{g})$ and $g(x;\theta_{g})$ respectively.
Methodology
In this section, we first reveal the issue of the increasing gap of features between natural and adversarial samples in AFT. Then we introduce a disentanglement-based method, to model and further remove the specific latent features leading to the feature gap. Our method is expected to mitigate the feature gap to enhance adversarial robustness.
Specific Latent Feature Causing Feature Gap
@@
As shown in Figure 1 (a), there is a growing gap in features between natural and adversarial samples in existing AFT methods. Intuitively, the ideal robust model should have a consistent understanding and analysis of natural and adversarial samples, thus extracting similar (the same) features from them. Thus, narrowing the feature gap is expected to achieve better robust performance.
@@
To realize the target, we can disentangle two types of features from the features of natural and adversarial samples. We define natural features $f_{nat}$ as the intrinsic features $f_{i}$. We can define adversarial features $f_{adv}$ as the combination of the intrinsic features $f_{i}$ and the confused features $f_{c}$, the latter of which leads to the gap in features between natural and adversarial samples. In this modeling, the intrinsic features represent the features that contribute to accurate classification, which are easily extracted from natural samples. Besides, the confused features are identified as the specific latent features causing the gap between features of natural and adversarial samples, which are confused by the adversarial perturbation. As illustrated in Figure 1 (b), if the confused features are removed in adversarial features, the features of adversarial samples can align with those of natural samples, resulting in a zero feature gap and accurate predictions for adversarial samples.
@@
We suppose the phenomenon in Figure 1 (a) is attributable to the presence of the confused features $f_{c}$. Because the confused features persist as latent features and don’t diminish during fine-tuning, the gap between features of natural and adversarial samples cannot be bridged. The confused features induce the gap between natural and adversarial features, leading to limited robust performance. Based on our observations and analyses, we introduce a new approach named Adversarial Fine-tuning via Disentanglement (AFD), as shown in Figure 2. It applies a feature disentangler to explicitly extract the confused features, and remove them by the elimination and alignment of features.
Adversarial Fine-tuning Based on Disentanglement
@@
The primary strategy is to explicitly disentangle the intrinsic and confused features from the features of adversarial samples. We propose a feature disentangler to model confused features for further elimination. The features in the last hidden layer are input into the disentangler, yielding two disentangled feature vectors. We define these feature vectors as intrinsic features $f_{1}$ and confused features $f_{2}$. To implement the disentanglement, we introduce a disentangler with two linear blocks $D_{1},D_{2}$ between the feature extractor $g$ and the linear classifier $\omega$. These linear blocks are expected to disentangle adversarial features into intrinsic and confused features. The disentangled feature vectors $f_{1}$ and $f_{2}$ are constrained to approximate the intrinsic and confused features. Given by:
@@
$\begin{split}&f_{1}={D_{1}}(g(x^{\prime};\theta_{g}^{F})),\\ &f_{2}={D_{2}}(g(x^{\prime};\theta_{g}^{F})),\\ &{\rm where}\ \left\{\begin{array}[]{lr}f_{1}\approx f_{i},&\\ f_{2}\approx f_{c},&\end{array}\right.\end{split}$ (1)
where $\theta_{g}^{F}$ denotes the parameters of the feature extractor in the fine-tuned model. $x^{\prime}$ denotes the adversarial samples created by PGD-10 (Madry et al. 2017), formulated as:
$x^{\prime}=\mathop{\boldsymbol{\Pi}}\limits_{\mathbb{N}(x,\epsilon)}\left(x+\epsilon\cdot\operatorname{sign}\left(\nabla_{x}\mathcal{L}_{CE}\left(\omega^{F}\cdot{D_{1}}(g(x;\theta_{g}^{F})),y\right)\right)\right),$ (2)
@@
where $\boldsymbol{\Pi}$ denotes a clamp function, $\operatorname{sign}$ denotes a sign function, $\mathcal{L}_{CE}$ denotes the cross-entropy loss. We take $f_{1}$ and $f_{2}$ to match the intrinsic and confused features. The input and output dimensions of $D_{1},D_{2}$ are set to match the channel dimensions of the feature vectors before disentanglement. This ensures the parameters of the linear block $\theta_{D_{1}}$ can be seamlessly integrated with the parameter of the last linear classifier $\omega^{F}$, expressed as $\theta_{D_{1}}\cdot\omega^{F}={\omega^{F\prime}}$. Therefore, we only need to load the linear classifier $\omega^{F\prime}$ without additional disentangling modules in the inference (test) phase.
@@
To achieve good performance on both accuracy and robustness, we should design suitable loss functions to realize Formula 1. We suggest fitting $f_{2}$ to incorrectly predicted labels $y^{\prime}$ to model the confused features. The confused features result in obvious shifts in both features and predicted probabilities, leading to wrong predicted class labels in most cases. Thus, it can be identified as the specific latent features causing the gap between natural and adversarial features. To decrease the dependence between the confused features and ground-truth labels, we force the disentangled features $f_{2}$ to match incorrect class labels, formulated as:
@@
$\begin{split}\mathcal{L}_{1}=\ &\mathcal{L}_{CE}(\omega^{F}\cdot f_{2},y^{\prime}),\\ {\rm where}\quad&y^{\prime}={\rm\mathop{max}\limits_{y^{\prime}\neq y}}\ (\omega^{F}\cdot f_{1}),\end{split}$ (3)
@@
where $y^{\prime}$ denotes wrong predicted labels with the maximum predicted probability. We only optimize the disentangling block $D_{2}$ instead of the entire model. This selective optimization prevents potential harm to the feature extractor that could result from learning the confused features.
@@
In addition, we introduce a constraint on $f_{1}$ to keep away from the confused features $f_{2}$, removing the confused features from the features of adversarial samples. The exclusive force between $f_{1}$ and $f_{2}$ encourages the elimination of the confused features from the features of adversarial samples. This constraint aims to protect adversarial features from being perturbed by the confused features, and reduces the association between confused and intrinsic features. Given by:
$\mathcal{L}_{2}=\ -\mathcal{D}(f_{1},f^{-}_{2}),$ (4)
@@
where $\mathcal{D}$ denotes a distance function, $f^{-}_{2}$ means the gradient back-propagation of $f_{2}$ is frozen.
Adversarial Fine-tuning Based on Alignment
@@
The second strategy is to align the adversarial features in the fine-tuned model with the natural features in the pre-trained model. Since the pre-trained model has a high natural accuracy and doesn’t suffer from adversarial attacks, the natural features in the pre-trained model can be identified as ideal intrinsic features. Thus, adversarial features in the fine-tuned model can be greatly improved by being aligned with the natural features in the pre-trained model. Besides, lots of work (Zhang et al. 2019; Wang et al. 2020; Suzuki et al. 2023) conducts an alignment between natural and adversarial features, which is regarded to mitigate the trade-off between accuracy and robustness. We take the disentangled features $f_{1}$ to match the natural features in the pre-trained model, further removing the confused features from $f_{1}$. The AFT loss function with the alignment can be formulated as:
$\mathcal{L}_{3}=\mathcal{L}_{CE}(\omega^{F}\cdot f_{1},y)+\gamma\cdot\mathcal{D}(f_{1},g(x;\theta_{g}^{P})^{-}),$ (5)
@@
where $\theta_{g}^{P}$ denotes parameters of the feature extractor in the pre-trained model, $\gamma$ denotes the weight of the alignment, and $g(x;\theta_{g}^{P})^{-}$ denotes the gradient back-propagation of $g(x;\theta_{g}^{P})$ is frozen. Therefore, the objective function of our AFD is defined as follows:
@@
where $\alpha$ and $\beta$ denote the weights of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The algorithm of AFD in Algorithm 1.